# Execution-Feedback Driven Test Generation from SWE Issues

### v1

@@ 6. Threats to Validity
One limitation of our work is that it only experiments with Python. As a result, the findings may not generalize to other programming languages or repositories outside the benchmarks. However, it is worth noting that prior works (Jimenez et al., 2024; Xia et al., 2025; Zhang et al., 2024), which share this same limitation, have still been impactful and contributed significantly to the field.
One of the key concerns with SWE-bench related work is the contamination or memorization problem of the model. The model may have seen the data during the training process, and that might positively impact the results. Prior works (Jimenez et al., 2024; Ahmed et al., 2025; Mündler et al., 2024) have attempted to address this problem. The SWE-bench paper claims that “difficulty does not correlate with issue resolution date”, indicating minimal contamination. Ahmed et al. show that Otter-generated tests differ substantially from developer-generated tests. In this work, we reuse Otter-generated tests. The only way we can introduce new contamination is while rewriting the issue description for morphs. However, Section 5.1 demonstrates through manual validation that issue morphs do not introduce contamination. We believe our results are not influenced by major model contamination.
Our system, e-Otter++, uses both sequential inference scaling (via test repair) and parallel inference scaling (via heterogeneous prompting). This increases the number of LLM calls. Furthermore, we also use Agentless code patches for test selection, which requires several LLM calls. We did not generate the code patches for Agentless ourselves; instead, we relied on code patches shared by the authors for both the Claude and GPT-4o models. However, based on their reported numbers and our experiments, we can estimate the cost for each sample. For the Claude and GPT-4o models, our estimated cost is \$2.75 and \$1.80 per instance, respectively. Note that prior test generation approaches do not use code patches for test selection, which adds more cost. However, it is likely that a reproduction test generation approach will be paired with a SWE-patch generation approach. In that case, the additional cost of generating code patches will be amortized.
Another limitation of e-Otter++ is that it considers only one test file and generates only one block of code. In real-world projects, test code can be spread across multiple files or blocks of code. Despite that, e-Otter++ exceeds the state-of-the-art performance, so we leave further improvements to future work.
@@ 7. Related Work
There has been a flurry of recent papers about generating reproduction tests from issues, which is the same problem we tackle in our paper. Libro generates tests for Java bug reports (Kang et al., 2023). It repairs tests based on compiler feedback, not execution feedback like we do, and generates several tests using high LLM decoding temperature, not morphs and masks like we do. Plein et al. also explore the feasibility of generating tests for Java bug reports, albeit without test repair (Plein et al., 2024). Mündler et al. (Mündler et al., 2024) introduce SWE-Agent+, a Python reproduction test generation agent derived from the SWE-Agent (Yang et al., 2024) code patch generator. While SWE-Agent+ has access to tools that can give it execution feedback, it leverages such feedback less effectively than e-Otter++. The same holds for EvoCoder (Lin et al., 2024), also derived from SWE-Agent, and for USEAgent (Applis et al., 2025), derived from AutoCodeRover (Zhang et al., 2024). AEGIS is a novel multi-agent system for reproduction test generation (Wang et al., 2024). It repairs tests based on execution feedback from $c_{\textrm{old}}$, but unlike our work, does not augment the test repair prompt. Also, it standardizes issue descriptions, but unlike our work, does not use morphs or masks to increase test candidate pool diversity. Otter++ (Ahmed et al., 2025) is an LLM-based workflow for reproduction test generation that we use as a component in e-Otter++, which extends it in three important ways: (i) augmenting the test repair prompt; (ii) using morphs (Otter++ only uses masks); and (iii) leveraging execution feedback from generated candidate code patches. Our experiments show that these novel contributions greatly improve the $F\!\!\to\!\!P$ rate. BRT Agent is a new agent for generating reproduction tests (Cheng et al., 2025), and Issue2Test is a new non-agentic workflow for the same task (Nashid et al., 2025); neither augments test repair prompts, uses masks or morphs, or leverages candidate code patches.
@@
Unlike e-Otter++ and the above papers on using generated code patches to help pick or improve reproduction tests, there is also work on the opposite: using reproduction tests to help pick or improve code patches. Papers on using tests to help pick (but not improve) code patches include Smith et al.’s 2015 paper (Smith et al., 2015), CodeT (Chen et al., 2023), Agentless (Xia et al., 2025), and R2E-Gym (Jain et al., 2025). Similarly, several papers use tests to help improve (but not pick) code patches (Zhang et al., 2024; Yang et al., 2024; Chen et al., 2024; Arora et al., 2024; Ruan et al., 2024; Mathews and Nagappan, 2024). Some papers even do both, using tests to help pick and improve code patches: CodeMonkeys (Ehrlich et al., 2025) runs several parallel workflows, each of which iteratively refines a $\langle c_{\mathrm{new}},y\rangle$ pair; and PatchPilot (Li et al., 2025) uses reproduction tests for bug localization before code patch generation, and later uses them again to pick among multiple candidate code patches. However, while most of the work described in this paragraph uses generated tests, test generation is not their main focus: their reproduction test generation components are more basic, and they lack empirical evaluation of the $F\!\!\to\!\!P$ rate of generated tests. That said, they demonstrate a widespread need for our work.

### v2

@@ 6. Threats to Validity
One limitation of our work is that it only experiments with Python. As a result, the findings may not generalize to other programming languages. However, it is worth noting that prior works (jimenez_et_al_2024; xia_et_al_2025; zhang_et_al_2024), which share this same limitation, have still been impactful and contributed significantly to the field. One of the key concerns with SWE-bench related work is the contamination or memorization problem of the model. To address this problem, we evaluated e-Otter++ on SWE-rebench and observed similar impact from execution-augmented test repair and heterogeneous prompting. Another limitation of e-Otter++ is that it considers only one test file and generates only one block of code. In real-world projects, test code can be spread across multiple files or blocks of code. Despite that, e-Otter++ exceeds the state-of-the-art performance, so we leave further improvements to future work.
It is hard to ensure the correctness of the generated tests. TDD-bench-Verified and SWT-bench use coverage in conjunction with the $F\!\!\to\!\!P$ property to be considered a success. Coverage is a primary metric for test quality. These benchmarks allow us to apply the golden patch (written by real developers and likely to be correct), run the test, and check both how well it covers the edited lines and whether it goes from fail to pass. While this is no correctness guarantee, it is a strong proxy for correctness and followed by prior works (mundler_et_al_2024; ahmed_et_al_2025). Besides, using imperfect patches or tests comes with certain risks, but prior work (chen_et_al_2023) has shown that dual execution agreement can help discard samples with potential errors. To further mitigate risk, we can use separate systems for surrogate patches and the proposed final patch (of course, with increased cost).
@@ 7. Related Work
There has been a flurry of recent papers about generating reproduction tests from issues, which is the same problem we tackle in our paper. Libro generates tests for Java bug reports (kang_yoon_yoo_2023). It repairs tests based on compiler feedback, not execution feedback like we do, and generates several tests using high LLM decoding temperature, not morphs and masks like we do. Plein et al. also explore the feasibility of generating tests for Java bug reports, albeit without test repair (plein_et_al_2024). Mündler et al. (mundler_et_al_2024) introduce SWE-Agent+, a Python reproduction test generation agent derived from the SWE-Agent (yang_et_al_2024) code patch generator. While SWE-Agent+ has access to tools that can give it execution feedback, it leverages such feedback less effectively than e-Otter++. The same holds for EvoCoder (lin_et_al_2024), also derived from SWE-Agent, and for USEAgent (applis_et_al_2025), derived from AutoCodeRover (zhang_et_al_2024). AEGIS is a novel multi-agent system for reproduction test generation (wang_et_al_2024). It repairs tests based on execution feedback from $c_{\textrm{old}}$, but unlike our work, does not augment the test repair prompt. Also, it standardizes issue descriptions, but unlike our work, does not use morphs or masks to increase test candidate pool diversity. Otter++ (ahmed_et_al_2025) is an LLM-based workflow for reproduction test generation that we use as a component in e-Otter++, which extends it in three important ways: (i) augmenting the test repair prompt; (ii) using morphs (Otter++ only uses masks); and (iii) leveraging execution feedback from generated candidate code patches. Our experiments show that these novel contributions greatly improve the $F\!\!\to\!\!P$ rate. BRT Agent is a new agent for generating reproduction tests (cheng_et_al_2025), and Issue2Test is a new non-agentic workflow for the same task (nashid_et_al_2025); neither augments test repair prompts, uses masks or morphs, or leverages candidate code patches.
@@
Unlike e-Otter++ and the above papers on using generated code patches to help pick or improve reproduction tests, there is also work on the opposite: using reproduction tests to help pick or improve code patches. Papers on using tests to help pick (but not improve) code patches include Smith et al.’s 2015 paper (smith_et_al_2015), CodeT (chen_et_al_2023), Agentless (xia_et_al_2025), and R2E-Gym (jain_et_al_2025). Similarly, several papers use tests to help improve (but not pick) code patches (zhang_et_al_2024; yang_et_al_2024; chen_et_al_2024; arora_et_al_2024; ruan_zhang_roychoudhury_2024; mathews_nagappan_2024). Some papers even do both, using tests to help pick and improve code patches: CodeMonkeys (ehrlich_et_al_2025) runs several parallel workflows, each of which iteratively refines a $\langle c_{\mathrm{new}},y\rangle$ pair; and PatchPilot (li_et_al_2025) uses reproduction tests for bug localization before code patch generation, and later uses them again to pick among multiple candidate code patches. However, while most of the work described in this paragraph uses generated tests, test generation is not their main focus: their reproduction test generation components are more basic, and they lack empirical evaluation of the $F\!\!\to\!\!P$ rate of generated tests. That said, they demonstrate a widespread need for our work.
@@ 8. Conclusion
Acknowledgements.
We gratefully acknowledge Ibragim Badertdinov and Chunqiu Steven Xia for sharing their model-generated SWE-rebench and Agentless patches.